Bwindi Impenetrable Forest (04.01.2014 to 06.30.2014)

You are here

PDF versionPDF version
April - June
Communications activities: 

 No presentations given in this period.

Unusual Events at the TEAM Site: 

 No unusual events at the site.

New Species: 

 No new species at the site.

Protocol Activities: 
Climate monitoring protocol
All climate data and maintenance logs were uploaded on to the TEAM portal. Bi-weekly maintenance of the climate station has been carried out in accordance with the protocol. 
Terrestrial vertebrate monitoring protocol
Array one (3o camera locations) deployment and retrieval was completed. Images from Array one have been downloaded in to DeskTEAM computer program. Camera trap setting for array 2 is in progress. 
Protocol Problems: 

 We are experiancing increased camera faillures. I suggest we purchase new cameras as soon as possible. We have hope to buy new camera traps for next year's sampling.

Schedule Problems: 

 No schedule problems. All activities were implemented on schedule.

Logistical Problems: 

 We request for more fund allocation towards equipment replacement. Our tree measuring diameter tapes are really faded. We need to replace the failling camera traps.


We need a forum/platform where site managers can share views/ideas on how to utilize our rapidly growing data base; potential data analysis and publications.